# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ B.2 Nonlinear problems
Appendix E Additional resources for the numerical experiments in Section 5
We begin with PDEs in strong form discretized via Physics-Informed Neural Networks (PINNs) [59, 19]. In all PINNs experiments, we set in Algorithm 1 the parameters $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=p$, $\mathtt{maxit}=20$, and $\kappa=0.1$. For NGD and NGD with CG, we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [72], from which most experiments are adapted. To enforce stronger regularization in early iterations, in NyströmNGD we take the regularization parameter as $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-4}\cdot L(\boldsymbol{\theta})^{2})$. Errors are always computed using an order of magnitude more quadrature points than in training, sampled independently. Random integration points sampled once and fixed throughout training.
E.1 Poisson problem in 3D
We consider the Poisson problem on the 3D unit cube
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ as in [72]. Problem (24) is linear and we consider it in strong form. Using the notation of Section B.1, the function spaces are therefore $\mathrm{V}=\mathrm{H}^{2}(\Omega)$ and $\mathrm{W}=\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\cong\mathrm{W}^{\prime}$, with operator
$\mathcal{A}:\mathrm{H}^{2}(\Omega)\to\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega),\qquad\mathcal{A}u=(-\Delta u,\left.{u}\right|_{\partial\Omega}).$
The corresponding least-squares loss is
and the induced metric reads
To discretize problem (24) using a PINN formulation, we take $u=u_{\boldsymbol{\theta}}$ a neural network and the loss function $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$ is approximated using MonteCarlo quadrature.
Remark E.1.
The metric (26) is coercive only in the $\mathrm{H}^{1/2}(\Omega)$ norm, but not in $\mathrm{H}^{2}(\Omega)$. This is due due to the use of the $\mathrm{L}^{2}(\partial\Omega)$ norm for the Dirichlet boundary term. Coercivity in $\mathrm{H}^{2}(\Omega)$ can be restored by enforcing the boundary conditions strongly [72, Remark 6], or by penalizing in the $\mathrm{H}^{1/2}(\partial\Omega)$ norm [13]. However, in line with standard PINNs practice, we retain the $\mathrm{L}^{2}(\partial\Omega)$ penalty for simplicity.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $500$ iterations and BFGS for $6\times 10^{4}$ iterations.
Figure 6: Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 7: Poisson equation in 3D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
E.2 Heat equation in 3+1D
We consider the heat equation
$\begin{cases}\partial_{t}u-\Delta u=f&I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\Omega,\end{cases}$ (27)
with exact solution $u(t,x,y,z)=\left(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\right)e^{-\pi^{2}t/4}$ as in [72]. We consider Equation 27 in strong form, with function spaces $\mathrm{V}=\mathrm{L}^{2}(I,\mathrm{H}^{2}(\Omega))\cap\mathrm{H}^{1}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{W}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))\times\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))\times\mathrm{L}^{2}(\Omega)$, and linear operator
$\mathcal{A}u=\big{(}\partial_{t}u-\Delta u,\;\left.{u}\right|_{I\times\partial\Omega},\;u(0,\cdot)\big{)}.$
The corresponding least squares loss is
with induced metric
Considerations on the coercivity of the metric similar to Remark E.1 apply here. For details, we refer to [72, Section 4.5]. To discretize problem (27) we take $u=u_{\boldsymbol{\theta}}$ a neural network and approximate $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$ using MonteCarlo quadrature.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{I\times\Omega}=3000$ points uniformly sampled in the interior $I\times\Omega$ and $N_{\Gamma}=800$ on the spatio-temporal boundary $\Gamma=(I\times\partial\Omega)\cup(\{0\}\times\Omega)$. We run NGD-based methods for $1000$ iterations and BFGS for $6\times 10^{4}$ iterations.
Figure 8: Heat equation in 3+1D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 9: Heat equation in 3+1D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
E.3 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [45], with Reynolds number $\mathrm{Re}=40$, following the setup in [38, Section 4.1]. The system of PDEs reads
with exact solution
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$, $\nu=\frac{1}{\mathrm{Re}}=\frac{1}{40}$, and $c$ is chosen such that $\int_{\Omega}p=0$. Following [38], we consider the function spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)\times\{p\in\mathrm{H}^{1}(\Omega):\int_{\Omega}p=0\}$ and $\mathrm{W}=\mathrm{L}^{2}(\Omega)^{2}\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\cong\mathrm{W}^{\prime}$. We write problem (30) in strong form as a nonlinear least squares problem with residual
$\mathcal{R}(u,p)=\begin{bmatrix}-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f\\ \nabla\cdot u\\ u-g_{D}\end{bmatrix}$ (31)
and corresponding energy functional
As a metric we consider the Gauss–Newton metric (9). We refer to [38] for a discussion on coercivity. To discretize (30), we take $(u,p)=(u_{\boldsymbol{\theta}_{1}},p_{\boldsymbol{\theta}_{2}})$ two neural networks with weights combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$, and approximate $L(\boldsymbol{\theta})=E(u_{{\boldsymbol{\theta}}_{1}},p_{{\boldsymbol{\theta}}_{2}})$ using Monte Carlo quadrature.
Computational details
We employ fully connected neural networks $u_{\boldsymbol{\theta}_{1}}$ and $p_{\boldsymbol{\theta}_{2}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations.
Figure 10: Kovasnay Flow discretized via PINNs. Top row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$. Bottom row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $p$ vs. iterations (first $1000$ iterations), and relative $\mathrm{H}^{1}(\Omega)$ error on $p$ vs. wall-clock time.
Figure 11: Kovasnay Flow discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
E.4 Source recovery problem for the Poisson equation in 3D
Given a penalization parameter $\eta>0$ and interior observations $u_{d}\in\mathrm{L}^{2}(\Omega)$, we consider the inverse problem of recovering both the solution $u$ and the source term $f$ from data by minimizing the Tikhonov-regularized objective:
$\inf_{(u,f)\in\mathrm{H}\times\mathrm{L}^{2}(\Omega)}J(u,f)=\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{\eta^{2}}{2}\left\|{f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (33)
subject to the Poisson PDE constraint
$\begin{cases}-\Delta u=f&\text{in }\Omega=(0,1)^{3},\\ u=g_{D}&\text{on }\partial\Omega,\end{cases}$ (34)
where $\mathrm{H}=\overline{C^{\infty}(\overline{\Omega})}^{\left\|{\cdot}\right\|_{\mathrm{H}}}$ with $\left\|{u}\right\|_{\mathrm{H}}=\left\|{\Delta u}\right\|_{\mathrm{L}^{2}(\Omega)}+\left\|{u}\right\|_{\mathrm{L}^{2}(\partial\Omega)}$. See [72, Section 4.7] for well-posedness and theoretical background. While the solution $u$ in [72] yields $f=0$, in this paper we choose a more complex problem, where $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ and the resulting source term to be reconstructed is not identically zero.
We discretize problem (3334) in strong form using a PINN formulation, where observations $u_{d}\in\mathrm{L}^{2}(\Omega)$ are in practice replaced by $N_{\mathrm{data}}$ point measurements. The function spaces are $\mathrm{V}=\mathrm{H}\times\mathrm{L}^{2}(\Omega)$, $\mathrm{W}=\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\mathrm{L}^{2}(\Omega)$, and the linear operator
$\mathcal{A}u=\left(\Delta u+f,u,\left.{u}\right|_{\partial\Omega},\eta f\right).$
The corresponding least squares loss is
$E(u,f)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}+\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{\eta^{2}}{2}\left\|{f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (35)
with induced metric
${g}\left({(u,f)},\,{(v,h)}\right)=\left({\Delta u+f},\,{\Delta v+h}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\Omega)}+\eta^{2}\left({f},\,{h}\right)_{\mathrm{L}^{2}(\Omega)}.$ (36)
We take $(u,f)=(u_{\boldsymbol{\theta}_{1}},f_{\boldsymbol{\theta}_{2}})$ two neural networks with weights combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$, and approximate $L(\boldsymbol{\theta})=E(u_{{\boldsymbol{\theta}}_{1}},f_{{\boldsymbol{\theta}}_{2}})$ using Monte Carlo quadrature and point observations of $u_{d}$. Although $\eta>0$ is theoretically required to ensure the minimum of (35) exists in $\mathrm{V}$, we set $\eta=0$ in our numerical experiments.
Computational details
We employ fully connected neural networks $u_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ with $\tanh$ activation and two hidden layers of width $32$. The term $\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ in the energy is discretized using $N_{\mathrm{data}}=40$ noiseless observations of the true solution at random points in $\Omega$. The remaining integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations.
Figure 12: Source recovery problem for the Poisson equation in 3D discretized via PINNs. Top row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$. Bottom row, from left to right: relative $\mathrm{L}^{2}(\Omega)$ error on $f$ vs. iterations (first $500$ iterations), and relative $\mathrm{L}^{2}(\Omega)$ error on $f$ vs. wall-clock time.
Figure 13: Source recovery problem for the Poisson equation in 3D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
Appendix F Additional Numerical Experiments: Variational Formulations of PDEs
We now consider variational formulations of PDEs, where the solution is approximated by a neural network $u_{\boldsymbol{\theta}}$. We consider formulations closely related to Variational PINNs (VPINNs) [41, 40] and the Deep Ritz method [25]. In all examples, we interpolate the neural network onto a conforming finite element (FE) space, following [7]. This approach offers several advantages. First, Dirichlet boundary conditions can be approximately enforced by applying them to the FE interpolant using standard FE techniques, avoiding the need to impose them directly on the neural network output [10, 9]. Second, it alleviates integration challenges, a known issue in neural network-based PDE solvers [61], especially in the Deep Ritz method [64]. Finally, spatial derivatives can be computed at the FE level, eliminating the need for automatic differentiation with respect to the input of the neural network. Further details are provided in [7, 6].
In all experiments involving variational formulations, we set the parameters in Algorithm 1 as follows: $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=30$, $\mathtt{maxit}=20$, and $\kappa=0.1$. The neural networks are interpolated on a finite element space of degree $d=4$ on a uniform quadrilateral mesh with mesh size $h=1/30$, yielding approximately $14,500$ degrees of freedom. Integrals in the loss and in the metric are evaluated by interpolation onto the FE space and Gauss-Jacobi quadrature of order $82d=$ 8. Errors are computed with a Gauss-Jacobi quadrature rule of order $164d=$ 16. We rely on the Julia packages Lux.jl [57] for neural networks, and Optim.jl [52] for optimization.
F.1 Poisson problem in 2D – Deep Ritz formulation
We consider the Poisson problem on the 2D unit square
The solution is then obtained as $u=u_{0}+\tilde{g}_{D}$. The spaces for the discretization are therefore $\mathrm{V}=\mathrm{W}=\mathrm{H}^{1}_{0}(\Omega)$, and the linear operator is given by
$\mathcal{A}:\mathrm{H}^{1}_{0}(\Omega)\to\mathrm{H}^{-1}(\Omega),\qquad\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}\mathbf{x}.$
Since $a$ is a symmetric and coercive bilinear form on $\mathrm{V}\times\mathrm{V}$, we can employ the Deep Ritz formulation and consider the minimization problem
$\begin{split}\min_{u_{0}\in\mathrm{H}^{1}_{0}(\Omega)}E(u_{0})&\coloneqq\int_{\Omega}\frac{1}{2}\left\|{\nabla u_{0}}\right\|^{2}-fu_{0}+\nabla\tilde{g}_{D}\cdot\nabla u_{0}\,\mathrm{d}\mathbf{x},\end{split}$ (39)
with corresponding metric is $g=a$.
To discretize (37), we approximate the solution $u=u_{0}+\tilde{g}_{D}$ as $u\approx\Pi u_{\boldsymbol{\theta}}+\tilde{g}_{D}$, where $u_{\boldsymbol{\theta}}$ is a neural network, $\Pi$ is the interpolation operator onto the non-Dirichlet degrees of freedom of the FE space, and an approximate lifting $\tilde{g}_{D}$ is obtained by interpolating $g_{D}$ on the boundary degrees of freedom using standard FE procedures. We then minimize the loss $L(\boldsymbol{\theta})=E(\Pi u_{\boldsymbol{\theta}})$.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations. Since the loss can become negative, for NGD and NGD with CG we cannot employ the heuristic $\mu=\min(L(\boldsymbol{\theta}),10^{-5})$ as suggested in [72], and therefore we set $\mu=10^{-9}$. In NyströmNGD we take $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-9})$.
Results
On this test case, NyströmNGD and BFGS achieve comparable final accuracy, as shown in Figure 14, but NyströmNGD requires a fraction of the iterations and is approximately 10×10$\times$10 × faster in wall-clock time. This efficiency arises from the excellent low-rank approximability of the Gramian, evident in Figure 15, which leads the Nyström rank parameter $\ell$ to stabilize around $80$, far below the parametric dimension $p$. NGD and NGD with CG are not competitive in either iteration count or runtime, and both plateau at lower accuracy.
Figure 14: Poisson equation in 2D discretized via Deep Ritz and finite element interpolation. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 15: Poisson equation in 2D discretized via Deep Ritz and finite element interpolation, trained with NGD using an SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
F.2 Diffusion-reaction-transport problem in 2D
We consider the diffusion-reaction-transport problem
The function spaces are $\mathrm{V}=\mathrm{W}=\mathrm{H}^{1}_{0,\Gamma_{D}}(\Omega)$, and the variational formulation of (40) is defined as $\mathcal{A}u=f$ where
$\begin{split}\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=a(u,v)=\int_{\Omega}(\kappa\nabla u)\cdot\nabla v+(\boldsymbol{\beta}\cdot\nabla u)\,v+\sigma\,u\,v\,\,\mathrm{d}\mathbf{x}\\ \left\langle{f},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=\int_{\Omega}F\,v\,\,\mathrm{d}\mathbf{x}+\int_{\Gamma_{N}}\psi\,v\,\,\mathrm{d}\mathbf{s}\end{split}$
Since $\mathcal{A}$ is in general not symmetric positive definite, we adopt the least-squares formulation
$\min_{u\in\mathrm{H}^{1}_{0}(\Omega)}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{H}^{-1}_{\Gamma_{D}}}^{2}=\left\langle{\mathcal{A}u-f},\,{\mathcal{R}_{\mathrm{V}}^{-1}(\mathcal{A}u-f)}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}},$ (41)
with associated metric ${g}\left({v},\,{w}\right)\coloneqq\left\langle{\mathcal{A}w},\,{\mathcal{R}_{\mathrm{V}}^{-1}\mathcal{A}v}\right\rangle$. Here, $\mathcal{R}_{\mathrm{V}}^{-1}=(-\Delta)^{-1}:\mathrm{H}^{-1}_{\Gamma_{D}}\to\mathrm{H}^{1}_{0,\Gamma_{D}}$ denotes the inverse Riesz operator, which involves solving a Poisson equation.
To discretize (41) and the corresponding metric, we interpolate the neural network $u_{\boldsymbol{\theta}}$ onto a conforming finite element space $\mathrm{V}_{h}\subset\mathrm{V}$, following [7], and minimize $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$. Applying the discrete Riesz operator $\mathcal{R}_{\mathrm{V}_{h}}^{-1}$ corresponds to solving a Poisson problem with FEM. To reduce computational cost, one could instead use a spectrally equivalent approximation, such as a few cycles of a geometric multigrid preconditioner. However, for simplicity, we solve the Poisson problem directly using FEM, computing a Cholesky factorization of the Laplacian matrix once.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations. For NGD and NGD with CG we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [72]. For NyströmNGD, to enforce stronger regularization in early iterations, we take $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-8}\cdot L(\boldsymbol{\theta})^{2})$.
Results
As illustrated in Figure 16, NyströmNGD reaches convergence in few iterations and achieves better accuracy compared to other NGD methods. The gap in convergence speed relative to BFGS is substantial in terms of both iterations and time. The smaller difference in wall-clock time between NyströmNGD and BFGS, compared to the previous variational example, can be attributed to the higher Nyström rank parameter $\ell$ at which NyströmNGD stabilizes. Although $\ell$ remains much lower than the parametric dimension $p$, it is higher than in earlier examples. This is due to the slower, yet still strong, decay of the singular values of the Gramian, as shown in Figure 17.
Figure 16: Diffusion-reaction-transport equation in 2D discretized via Finite Element Interpolated Neural Networks. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 17: Diffusion-reaction-transport equation in 2D discretized via Finite Element Interpolated Neural Networks, trained with NGD using an SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.

### v3

@@ A.2 Nonlinear problems
Appendix C Sensitivity analysis and hyperparameter tuning
We conduct a sensitivity analysis of NyströmNGD with respect to its key hyperparameters, using the 3D Poisson problem discretized via PINNs as a model problem. Specifically, we examine the impact of the regularization parameter $\mu$, the adaptivity of the Nyström rank parameter $\ell$, and the maximum number of iterations and tolerance for the CG algorithm. Notation follows Section˜4.2.
Regularization parameter $\mu$
We set $\mu=\gamma\cdot\epsilon_{\textrm{mach}}\cdot\hat{\lambda}_{1}$ to regularize the least squares problem, solving with $(\mathbf{G}(\boldsymbol{\theta})+\mu I)^{-1}$ instead of $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$. We recommend $\gamma=p$, following the threshold for numerical rank used in standard software like NumPy and MATLAB. Setting $\gamma$ much larger than $p$ leads to excessive regularization and a search direction that deviates from the “exact” natural gradient, while setting it much smaller can introduce numerical instability. This behavior is confirmed by the results in the Table˜9. Also, lowering $\gamma$ increases the Nyström rank parameter $\ell$, raising computational cost, as reflected in the relative iteration times.
Table 9: Sensitivity analysis of the parameter $\gamma$. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that with the default value $\gamma=p$.
$\boldsymbol{\gamma}$ Final accuracy Time (relative)
$\boldsymbol{p\times 100}$ $1.52\text{\times}{10}^{-04}$ $0.8$
$\boldsymbol{p\times 10}$ $9.21\text{\times}{10}^{-05}$ $0.9$
$\boldsymbol{p}$ $4.01\text{\times}{10}^{-05}$ $1.0$
$\boldsymbol{p\times 0.1}$ $70\text{\times}{10}^{-05} 1.1$1.1
$\boldsymbol{p\times 0.01}$ $3.91\text{\times}{10}^{-03}$ $1.3$
Rank adaptivity
We adapt the Nyström rank parameter $\ell$ as the smallest index satisfying $\hat{\lambda}_{\ell}/\mu<r$, following the recommendation $r=10$ from [31]. Decreasing $r$ increases the rank $\ell$, improving the quality of the preconditioner at the cost of higher computational expense. This trade-off is clearly reflected in the results in Table˜10. We find that the default value $r=10$ achieves a good balance between accuracy and per-iteration efficiency.
Table 10: Sensitivity analysis of the parameter $r$. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that with the default value $r=10$.
$\boldsymbol{r}$ Final accuracy Time (relative)
1000 $2.67\text{\times}{10}^{-04}$ $0.8$
100 $1.18\text{\times}{10}^{-04}$ $0.9$
10 $4.01\text{\times}{10}^{-05}$ $1.0$
1 $2.21\text{\times}{10}^{-05}$ $1.2$
0.1 $1.89\text{\times}{10}^{-05}$ $1.3$
Conjugate Gradient
We set the CG relative tolerance as $\mathtt{rel\_tol}=\min(\kappa,\lVert\nabla L(\theta)\rVert_{2})$ following the Trust-Region method, which preserves the local quadratic convergence of Newton’s method while also ensuring global convergence and reducing computation far from critical points [20]. To obtain steps closer to the “exact” NGD direction, one can decrease $\kappa$ and increase the maximum number of CG iterations $\mathtt{maxit}$. As shown in Table˜11, increasing $\mathtt{maxit}$ and lowering $\kappa$ slightly raises runtimes, though the main per-iteration cost remains dominated by computing the preconditioner. Final accuracy varies minimally, and the default values $\kappa=0.1$ and $\mathtt{maxit}=20$ provide a good balance between accuracy and efficiency.
Table 11: Sensitivity analysis of the CG parameters $\kappa$ and $\mathtt{maxit}$. Final accuracy (relative $\mathrm{H}^{1}$ error), and in parentheses the iteration time relative to that with the default values $\kappa=0.1$ and $\mathtt{maxit}=20$.
$\boldsymbol{\kappa}$ $\boldsymbol{\mathtt{maxit}}$ 2 10 20 50 100
$\boldsymbol{10^{-1}}$ $3.72\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$)
$\boldsymbol{10^{-2}}$ $3.52\text{\times}{10}^{-05}$ ($1.1$) $4.12\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-3}}$ $4.59\text{\times}{10}^{-05}$ ($1.1$) $5.65\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-6}}$ $3.33\text{\times}{10}^{-05}$ ($1.0$) $1.48\text{\times}{10}^{-05}$ ($1.1$) $1.70\text{\times}{10}^{-05}$ ($1.1$) $1.85\text{\times}{10}^{-05}$ ($1.1$) $1.78\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-10}}$ $3.33\text{\times}{10}^{-05}$ ($1.0$) $1.43\text{\times}{10}^{-05}$ ($1.1$) $1.52\text{\times}{10}^{-05}$ ($1.1$) $2.00\text{\times}{10}^{-05}$ ($1.3$) $1.98\text{\times}{10}^{-05}$ ($1.3$)
Appendix D Alternative preconditioners
The matrix $(\mathbf{G}(\boldsymbol{\theta})+\mu I)$ that we aim to precondition in NGD is symmetric positive definite, dense, well-approximated by a low-rank matrix, and accessible efficiently through matrix-vector products, but not in $O(1)$ elementwise time. In this setting, to the best of our knowledge, randomized Nyström preconditioning represents the state of the art for constructing an efficient preconditioner. To address concerns about the suitability of the Nyström approximation, we also tested alternative low-rank preconditioners, namely Pivoted Partial Cholesky, with greedy pivot selection [43], uniformly random pivot selection, and RPCholesky pivot selection [17]. Results are reported in Table˜12.
As already noted in Remark˜4.5, the main limitation of greedy pivot selection and RPCholesky in our context is that they require access to the entire diagonal of $\mathbf{G}(\boldsymbol{\theta})$, which involves computing matvecs with all canonical basis vectors. This is reflected in the longer runtimes in Table˜12 and would be exacerbated for larger networks, whereas NyströmNGD scales efficiently, as shown in Section˜5.1.2. In contrast, selecting pivots uniformly at random produces a significantly less effective preconditioner. Moreover, as noted in [17, Property 2.1], Pivoted Partial Cholesky is mathematically equivalent to Nyström with sampling matrices corresponding to the pivots. This supports the use of Gaussian random vectors in Nyström in our setting: they achieve similar accuracy without requiring explicit pivot selection.
Table 12: Comparison of different preconditioners for matrix-free NGD. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that of Nyström NGD.
Preconditioner Final accuracy Time (relative)
NyströmNGD $2.92\text{\times}{10}^{-05}$ 11
Greedy Pivoted Chol. $2.18\text{\times}{10}^{-05}$ $4.5$
RPCholesky $2.61\text{\times}{10}^{-05}$ $4.7$
Randomly Pivoted Chol. $1.22\text{\times}{10}^{-02}$ $0.62$